### v1

# Weakly Supervised Gaussian Contrastive Grounding with Large Multimodal Models for Video Question Answering

Abstract
Video Question Answering (VideoQA) aims to answer natural language questions based on the information observed in videos. Despite the recent success of Large Multimodal Models (LMMs) in image-language understanding and reasoning, they deal with VideoQA insufficiently by simply taking uniformly sampled frames as visual inputs, which ignores question-relevant visual clues. Moreover, there are no human annotations for question-critical timestamps in existing VideoQA datasets. In light of this, we propose a novel weakly supervised framework to enforce the LMMs to reason out the answers with question-critical moments as visual inputs. Specifically, we fuse the question and answer pairs as event descriptions to find multiple keyframes as target moments, which will be pseudo-labels. With these pseudo-labels as additionally weak supervision, we devise a lightweight Gaussian-based Contrastive Grounding (GCG) module. GCG learns multiple Gaussian functions to characterize the temporal structure of the video, and sample question-critical frames as positive moments to be the visual inputs of LMMs. Extensive experiments on several VideoQA benchmarks verify the effectiveness of our framework, and we achieve substantial improvements compared to previous state-of-the-art methods
✉ Corresponding AuthorCodes will be released after review
Figure 1: We argue that the information in uniformly sampled frames is insufficient for LMMs to answer the question correctly. Therefore, we utilize the fused event description to provide additional weak supervision and generate weight distributions for each video moment. We align the positive description-moment pairs while pushing away negative ones. In this way, we can effectively select question-critical moments for LMMs to reason out the answer.
1 Introduction
Video Question Answering (VideoQA) stands at the forefront of developing intelligent systems that can reason about causal and temporal relations and answer natural language questions in videos, which is an essential manifestation of human intelligence. Despite significant advancements in recent years, VideoQA remains a challenging problem that requires models to comprehensively understand and dynamically align the semantics of both the visual world and natural language.
With the notable progress in vision-language pre-training techniques Radford et al. (2021); Li et al. (2022b) and Large Language Models (LLMs) Touvron et al. (2023); Brown et al. (2020), Large Multimodal Models (LMMs) Li et al. (2023c); Dai et al. (2023); Liu et al. (2023) have showcased impressive capabilities across various image-language tasks. However, limited by the long sequence frames in videos and computation costs, current LMMs fall short when applied to VideoQA. They simply concatenate the visual tokens of uniformly/randomly sampled, sparse frames (e.g., 4 frames) as the visual inputs for answer prediction. Such a uniform sampling strategy does not consider the specific question, treating all frames equally and introducing redundancy, particularly when tackling complicated questions within longer videos.
Therefore, it’s necessary to localize the moments crucial for answering the question for LMMs (as the positive moments shown in Figure 1). Notably, different from the task of Temporal Sentence Grounding (TSG) Gao et al. (2017); Zhang et al. (2023) which aims to localize a video moment described by a declarative sentence, the grounding mechanism in VideoQA features some unique challenges. First, questions in VideoQA are interrogative sentences, and they lack explicit information about the answer content needed to be grounded. For instance, there is a semantic gap between the interrogative question [How does the boy react after opening the present?] and the declarative description [The boy carries up the dog after opening the present.]. Thus, models are required to both localize the moment [after opening the present] explicitly shown in the question and identify the implicit answer moment [carry up the dog], demanding the causal-temporal reasoning. Second, VideoQA aims to correctly answer the questions of videos, rather than solely grounding specific video moments, and there are no human annotations for the timestamps of question-critical moments in existing VideoQA datasets.
To address these challenges, we introduce a weakly supervised framework by discovering question-critical moments with Gaussian-based Contrastive Grounding (GCG). As labeling the timestamps of question-critical moments is labor-intensive and subjective, we leverage the visual-language alignment capability of the CLIP models Fang et al. (2023) to weakly provide timestamps of question-critical moments. In detail, we fuse the textual question and answer to generate a declarative sentence as the event description, and then compute the similarities between the description and each frame. Frames with the highest scores will be the keyframes of target moments. We observe that LMMs with these keyframes as visual inputs showcased significant improvements in VideoQA datasets. With these pseudo-labeled timestamps as additional weak supervision, our GCG will learn multiple Gaussian functions to characterize the inherent temporal structure of the video, and distinguish the positive video moments (green in Figure 1) from negative video moments (orange and yellow in Figure 1). The positive moments are crucial for answering the question and will be the visual inputs of LMMs for final answer prediction. Moreover, to ensure the selected positive moments are closest to the event description, our GCG also includes a contrastive objective He et al. (2020) that can align the positive description-moment pairs while pushing away negative ones. Different from previous works like SeViLA Yu et al. (2023) which pre-train an additional LMM as the keyframe localizer with other TSG datasets like QV-Highlights Lei et al. (2021), our method is lightweight for end-to-end training and flexible for LMMs.
To summarize, we make the following contributions:
  • We propose a weakly supervised grounding framework for VideoQA, by utilizing the alignment capability of the CLIP models to provide pseudo-labeled timestamps of keyframes without human-labor annotated costs.
  • We devise the Gaussian-based Contrastive Grounding (GCG) for weakly-grounded selection of question-critical moments, enhancing the efficiency and interpretability of LMMs when applied to VideoQA.
  • We conduct extensive experiments to verify the effectiveness of our proposed method, and achieve substantial improvements on six challenging VideoQA benchmarks including NExT-QA, Causal-VidQA, Intent-QA, ActivityNet-QA, MSVD-QA, and MSRVTT-QA.
2 Related Works
2.1 Large Multimodal Models (LMMs)
LMMs Dai et al. (2023); Li et al. (2023c); Ye et al. (2023); Liu et al. (2023); Alayrac et al. (2022) in their current form primarily function as image-to-text generative models, taking images as input and generating text sequences. These models have demonstrated strong capabilities in image-language understanding and reasoning by adapting frozen language models to frozen image encoders with trainable connection modules, following large-scale image-text pretraining. The connection module can either be a transformer-based architecture like Q-former in InstructBLIP and BLIP-2 Dai et al. (2023); Li et al. (2023c), Perceiver Resampler in Flamingo Alayrac et al. (2022), or a simple linear layer in LLaVA Liu et al. (2023). Most current LMMs are essentially image-based models, and they simply concatenate the visual tokens extracted from uniformly sampled, sparse frames as visual inputs for video-language tasks. This results in a lack of temporal modeling ability and emphasizes the necessity of selecting specific video moments, particularly for addressing the demands of reasoning-based VideoQA tasks. In this paper, our goal is to enhance the causal-temporal reasoning abilities of LMMs without additional pretraining on video-text corpora, by discovering the question-critical moments with our weakly-supervised Gaussian-based Contrastive Grounding.
2.2 Temporal Grounding in VideoQA
Early VideoQA benchmarks Yu et al. (2019); Xu et al. (2017); Jang et al. (2017) focus on descriptive questions (e.g., [what’s the man doing]) within short video clips, rarely going beyond a recognition of the objects and actions. Instead, more recent VideoQA benchmarks Li et al. (2022a); Xiao et al. (2021); Li et al. (2023b) like NExT-QA emphasize counterfactual, temporal, and causal reasoning involving multiple entities and relations, demanding the ability to uncover the causes behind specific events within longer videos, necessitating localizing a text query to specific moments.
In light of this, ATP Buch et al. (2022) utilizes the tool of atemporal probe to select a single frame without temporal information for downstream tasks. MIST Gao et al. (2023) and TranSTR Li et al. (2023e) fuse frames with the mechanism of adaptive temporal rationalization and iterative spatial-temporal attention, respectively. NExT-GQA Xiao et al. (2023) constructs grounding labels in the test set of NExT-QA and uses a single Gaussian mask to fuse frames along the temporal dimension. SeViLA Yu et al. (2023), similar to us, utilizes the LMM (BLIP-2) for VideoQA. However, SeViLA uses two LMMs to generate pseudo-labels and answer questions respectively, with extra pre-training on TSG datasets Lei et al. (2021) and a multi-stage training scheme. Different from previous works, we utilize the CLIP Radford et al. (2021); Fang et al. (2023) models to automatically provide weak supervision for grounding, and our lightweight GCG module learns multiple Gaussian functions to generate both positive and negative moments in an end-to-end manner, with an additionally contrastive objective to distinguish positive ones from negative ones for frame selection.
3 Preliminary: LMMs for VideoQA
We take InstructBLIP Dai et al. (2023) as an example to illustrate how LMMs deal with VideoQA. LMMs approach VideoQA as a text generation task conditioned on the question $\mathcal{Q}$ and video $\mathcal{V}$ with $T$ frames. Similar to other LMMs, InstructBLIP predicts the answer $\mathcal{A}$ by the following steps:
(1) The Vision Transformer Dosovitskiy et al. (2021) in EVA-CLIP Fang et al. (2023) serves as the frozen image encoder to extract embeddings of each frame individually, and obtains $\textbf{E}=\left\{\textbf{e}_{1},\textbf{e}_{2},\cdots,\textbf{e}_{T}\right\},\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}},\textbf{e}_{t}\in\mathbb{R}^{N_{I}\times D_{I}}$, where $t$ denotes the $t$-th frame, $N_{I}$ is the patch number of each frame (including the class token), and $D_{I}$ is the embedding dimension. To mitigate computational costs, existing LMMs uniformly sample $K$ frames ($K\ll T$) to represent the video, resulting in the $\hat{\textbf{E}}\in\mathbb{R}^{K\times N_{I}\times D_{I}}$.
(2) A trainable Q-former serves as the connection module to bridge the modality gap between vision and language. It takes frame embeddings $\hat{\textbf{E}}$ and outputs a set of fixed-length frame tokens $\textbf{F}=\left\{\textbf{f}_{1},\textbf{f}_{2},\cdots,\textbf{f}_{K}\right\},\textbf{F}\in\mathbb{R}^{K\times N_{C}\times D_{C}},\textbf{f}_{t}\in\mathbb{R}^{N_{C}\times D_{C}}$, where $N_{C}$ is the token number of each frame ($N_{C}\ll N_{I}$, e.g., $N_{C}=32$ and $N_{I}=257$ in InstructBLIP), and $D_{C}$ is the dimension of the connection module.
(3) Each $\textbf{f}_{t}$ in F are concatenated together to obtain the flattened $\textbf{F}\in\mathbb{R}^{(K\cdot N_{C})\times D_{C}}$, followed with a fully-connected layer to project F into the LLM’s dimension $D_{L}$. At last, the final projected $\textbf{F}\in\mathbb{R}^{(K\cdot N_{C})\times D_{L}}$ is fed into the frozen LLM (e.g., FLAN-T5 Chung et al. (2022) or Vicuna Zheng et al. (2023)) serving as soft prompts, together with the word embeddings of question $\mathcal{Q}$, to generate the answer text $\mathcal{A}$.
The model is trained by optimizing the trainable parameters $\theta$ of the model $P$ with the autoregressive objective:
$\mathcal{L}_{vqa}=-\sum_{t=1}^{L_{a}}logP_{\theta}(\mathcal{A}_{t}|\mathcal{A}_{<t},\mathcal{V},\mathcal{Q})$ (1)
where $\mathcal{A}_{t}$ is predicted autoregressively at position $t$, and $L_{a}$ is the sequence length of the ground truth answer text $\mathcal{A}$. Our motivation is to replace the uniformly sampled frames $\hat{\textbf{E}}$ in step (1) with question-critical frames as visual inputs.
Figure 2: (a) The overall framework of our method. (b) The process of pseudo-label generation.
4 Method
Figure 2 (a) gives an overview of our framework. After extracting the embeddings of $T$ frames, our GCG module will select the most question-critical $K$ frames $\hat{\textbf{E}}$ from E, as the visual inputs for the LMM. To ensure the selected frames are most relevant to answering the question, the GCG module will be additionally optimized by the pseudo-labels of question-critical moments, resulting in the regression objective $\mathcal{L}_{reg}$ from weakly grounded timestamps, and the contrastive objective $\mathcal{L}_{con}$ aligning the paired description-moment pairs while pushing away unpaired ones.
4.1 Inputs Representation
For video representations, along with the frame embeddings $\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}}$ extracted in step (1) of Section 3, we also extract the corresponding class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ from E for further pseudo-labels generation and contrastive grounding. For language representations, we tokenize the question $\mathcal{Q}$ into a sequence of words and then feed them into the text encoder of EVA-CLIP, to get word-level embeddings $\textbf{Q}=\left\{\textbf{q}_{t}\right\}_{t=1}^{L_{q}}\in\mathbb{R}^{L_{q}\times D_{I}}$, where $L_{q}$ denotes the sequence length of the question. We get the embeddings of the fused event description (detailed in 4.2) the same way as Q, but only retain the class token $\textbf{d}_{[CLS]}\in\mathbb{R}^{D_{I}}$ to represent it for further pseudo-labels generation.
Figure 3: We use the Gaussian generator to generate multiple Gaussian functions and obtain weight distributions $p\in\mathbb{R}^{T}$ for each video moment. The Gaussian generator will be optimized by the weakly-supervised regression objective $\mathcal{L}_{reg}$ and contrastive objective $\mathcal{L}_{con}$, along with the fully supervised QA objective $\mathcal{L}_{vqa}$, to discover the most question-critical moments as visual inputs for LMMs.
4.2 Pseudo Labels for Temporal Grounding
Considering the powerful visual-language alignment ability of EVA-CLIP, as in Figure 2 (b), we utilize its joint-trained image and text encoder to provide pseudo labels for timestamps of question-critical moments as weak supervision.
Event Description Generation. To adapt the textual representation for better event description and reduce the semantic gaps, we prompt the LLaMA Touvron et al. (2023) (or any other open-sourced LLM) to fuse the question and answer pairs with hand-written demonstrations. For example, the QA-pair [Q: How does the boy react after opening the present? A: carry up the dog] will be transformed into the declarative event description [The boy carries up the dog after opening the present.]. The event descriptions provide more accurate textual descriptions for question-critical moments because the answer content is included.
Pseudo Labels Generation. As in Figure 2 (b), we represent the video and event description with the class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ and $\textbf{d}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ respectively, to obtain the weakly labeled question-critical timestamps. In detail, we compute the cosine similarities between $\textbf{E}_{[CLS]}$ and $\textbf{d}_{[CLS]}$ and get the similarities scores $s\in\mathbb{R}^{T}$, which recording the relevance between each frame and the event description. Then, we choose the indexes of the highest Top-$K$ scores in $s$ as $w\in\mathbb{R}^{K}$ where $w_{k}\in\left\{1,2,\cdots,T\right\}$, to be the timestamps of question-critical frames. We verify the effectiveness of the weakly grounded $w$ in the analysis 5.4.
4.3 Gaussian Generator
Recent research Xiao et al. (2023) highlights the superiority of end-to-end Gaussian mask learning for grounding in VideoQA. Motivated by this, we design the effective weakly supervised Gaussian-based Contrastive Grounding (GCG). However, unlike Xiao et al. (2023) to generate a single Gaussian function, we generate multiple Gaussian functions to characterize the multi-event temporal structure of the video. Moreover, our GCG is optimized from both the QA supervision $\mathcal{L}_{vqa}$ and weakly labeled supervision, which is more effective for discovering the question-critical moments.
Specifically, we utilize the Gaussian generator to obtain $K$ Gaussian functions $g=\left\{g_{1},\cdots,g_{K}\right\},g_{k}\in\mathbb{R}^{T}$, depending on the video and question. These Gaussian functions will be combined into an overall weight distribution $p\in\mathbb{R}^{T}$ for each video moment. Notably, $p$ tends to have $K$ peaks, representing the most question-critical $K$ frames, with corresponding indexes to be the centers of each Gaussian function.
As in the right part of Figure 3, the Gaussian generator consists of a cross-modal embedding layer and a transformer encoder Vaswani et al. (2017). The cross-modal embedding layer is a down-sampling linear layer with learnable modal-type embeddings and positional embeddings. The concatenated multimodal embeddings $\textbf{M}=[\textbf{E}_{[CLS]};\textbf{Q}]\in\mathbb{R}^{(T+L_{q})\times D_{I}}$ serve as inputs for the Gaussian generator:
M $=\operatorname{Linear}(\textbf{M}),\ \textbf{M}\in\mathbb{R}^{(T+L_{q})\times D_{G}}$ (2)
$\textbf{M}[:T]=\textbf{M}[:T]\mathrel{+}\mathrm{Type}_{V}\mathrel{+}\mathrm{Pos};$
$\textbf{M}[T:]=\textbf{M}[T:]\mathrel{+}\mathrm{Type}_{T}$
Next, a standard transformer encoder is adopted to establish the cross-frame dynamics and cross-modal interactions, which takes the embedded M and yields $\hat{\textbf{M}}\in\mathbb{R}^{T\times D_{G}}$ (only reserve the first $T$ embeddings). Then, we use attention pooling to summarize the outputs $\hat{\textbf{M}}$ along the temporal dimension and derive the global video representations $\textbf{G}\in\mathbb{R}^{D_{G}}$. As G integrates all the video and question information, we predict the centers $\mu\in\mathbb{R}^{K}$ of $K$ learnable Gaussian functions weighting over the entire video sequence, through G with a fully connected head activated by Sigmoid function:
$\mu=\operatorname{Sigmoid}(\operatorname{Linear}(\textbf{G})),\ \mu\in\mathbb{R}^{K}$ (3)
Given predicted $\mu$, we get $K$ Gaussian functions $g=\left\{g_{1},\cdots,g_{K}\right\}$ parameterized with ($\mu$, $\sigma$):
$g_{k}$ $=\frac{1}{\sqrt{2\pi}\sigma}\mathrm{exp}(-\frac{(t/T-\mu_{k})^{2}}{2\sigma^{2}}),\ g_{k}\in\mathbb{R}^{T}$ (4)
$k=\left\{1,2,\cdots,K\right\},\ t=\left\{1,2,\cdots,T\right\}$
where $\sigma$ is a hyperparameter controlling the width of the Gaussian curve. Then, the weight distribution $p\in\mathbb{R}^{T}$ of each video moment is generated by summing each $g_{k}$:
$p=\mathrm{Norm}(\sum_{k=1}^{K}g_{k}),\ p\in\mathbb{R}^{T}$ (5)
$\mathrm{Norm}$(·) scales values into the range [0, 1]. As the $K$ peaks in $p$, whose corresponding indexes are $\left\{\mu_{1},\cdots,\mu_{K}\right\}$, represent the most question-critical $K$ frames, we optimize the Gaussian generator with the regression objective to measure the discrepancy between the predicted centers $\mu\in\mathbb{R}^{K}$ and the weakly grounded timestamps $w\in\mathbb{R}^{K}$ by smooth $L1$ loss:
$\mathcal{L}_{reg}=\sum_{k=1}^{K}\mathrm{Smooth}_{L1}\|\mu_{k}-w_{k}/T\|$ (6)
4.4 Contrastive Grounding
Contrastive grounding aims to ensure the selected moments are most relevant to the event description. To achieve this, we learn a cross-modal embedding space, where the embeddings of the event description $\textbf{d}_{[CLS]}$ should be well aligned with the selected positive frames $\textbf{E}^{pos}\in\mathbb{R}^{K\times D_{I}}$, which are derived from the weight distribution $p$, and far away from those irrelevant ones. $\textbf{E}^{pos}$ are also the class tokens of the selected frame embeddings $\hat{\textbf{E}}\in\mathbb{R}^{K\times N_{I}\times D_{I}}$, which will be the visual inputs of LMMs for final answer prediction.
Positive Moments Selection. Since the distribution $p$ weights each video moment based on its contribution to the question, we select the Top-$K$ elements from $\textbf{E}_{[CLS]}$ according to $p$, and obtain $\textbf{E}^{pos}\in\mathbb{R}^{K\times D_{I}}$ as the positive frames. However, the selection via vanilla hard Top-$K$ produces a discrete selection, making it inapplicable for end-to-end training. We address this issue by adopting a differentiable Top-$K$ using the perturbed maximum method Berthet et al. (2020).
Negative Moments Mining. To distinguish highly confusing scenes, we mine negative moments within the same video as intra-negative frames $\textbf{E}^{intra}\in\mathbb{R}^{N_{intra}\times D_{I}}$, by sampling frames with the lowest $N_{intra}$ weights in $p$ from $\textbf{E}_{[CLS]}$. We also use $N_{inter}$ frames randomly sampled from other videos within the same batch to serve as inter-negative frames $\textbf{E}^{inter}\in\mathbb{R}^{N_{intra}\times D_{I}}$. These negative samples from both the same video and other videos can provide richer information. The objective of contrastive grounding is described as:
$\mathcal{L}_{con}=-\frac{1}{K}\sum_{k=1}^{K}\mathrm{log}\frac{\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{pos}_{k}/\tau)}{\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{pos}_{k}/\tau)+\mathrm{SUM}}$ (7)
$\mathrm{SUM}=\sum_{i=1}^{N_{intra}}\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{intra}_{i}/\tau)\ +$
$\sum_{j=1}^{N_{inter}}\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{inter}_{j}/\tau)$
where $\tau$ is the temperature factor and $\otimes$ is the dot product. Contrastive grounding can maximize the similarity between the query $\textbf{d}_{[CLS]}$ and a group of corresponding positive video moments $\textbf{E}^{pos}$ under the joint embedding space while pushing away negative video moments.
4.5 Answer Prediction
With the distribution $p$ optimized by both $\mathcal{L}_{reg}$ and $\mathcal{L}_{con}$, we select the most weighted $K$ frame embeddings from $\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}}$ based on the $p$, and obtain the selected frame embeddings $\hat{\textbf{E}}\in\mathbb{R}^{K\times N_{I}\times D_{I}}$. This process replaces the uniform sampling, and the same perturbed maximum method is adopted for differentiability. At last, we feed $\hat{\textbf{E}}$ into the Q-Former and LLM as the steps (2) and (3) in Section 3 to autoregressively predict the answer $\mathcal{A}$. During training, the whole pipeline is optimized by the joint objective:
$\mathcal{L}=\mathcal{L}_{vqa}+\alpha_{1}\mathcal{L}_{reg}+\alpha_{2}\mathcal{L}_{con}$ (8)
where $\alpha_{1}$ and $\alpha_{2}$ are the hyper-parameters to control the strengths of GCG. During the inference process, GCG only generates the weight distribution $p$ with the Gaussian generator as in Section 4.3, and obtains the most weighted $K$ frame embeddings $\hat{\textbf{E}}$ via a fully discrete Top-$K$ selection.
5 Experiments
5.1 Datasets
NExT-QA Xiao et al. (2021) contains 5.4k videos with an average length of 44s and 52k QA pairs, including question types of description, causal, and temporal. Intent-QA Li et al. (2023b) focuses on intent reasoning in daily social activities, with more than 4.3k videos and 16k QA pairs, including question types of causal-why, causal-how, and temporal. Causal-VidQA Li et al. (2022a) selects 27k video clips and asks 108k questions, including types of description, explanation, prediction, and counterfactual. MSVD-QA and MSRVTT-QA Xu et al. (2017) emphasize the description of video objects, activities, and their attributes, with 50k QA pairs over 1,970 videos and 243K QA pairs over 10K videos respectively. ActivityNet-QA Yu et al. (2019) consists of 58k QA pairs on 5.8k long web videos, with an average length of 180 seconds. NExT-QA, Intent-QA, and Causal-VidQA use a multi-choice setting to test temporal reasoning with causal and commonsense relations. MSVD-QA, MSRVTT-QA, and ActivityNet-QA employ an open-ended setting, focusing on the description of different elements in videos.
5.2 Implementation Details
We choose InstructBLIP Dai et al. (2023) as our LMM, with EVA-CLIP Fang et al. (2023) as the image encoder and text encoder, Q-former Li et al. (2023c) as the connection module, and Vicuna Zheng et al. (2023) or FLAN-T5 Chung et al. (2022) as the large language model. We sample each video as a sequence of $T=32$ frames and select $K=4$ frames as visual inputs. The number of negative samples is $N_{intra}=16$ and $N_{inter}=32$. The number of transformer encoder layers in the Gaussian generator is 2, with the hidden size $D_{G}=256$. For the hyperparameters, we set $\sigma=0.2$, $\tau=0.1$, $\alpha_{1}=\alpha_{2}=0.1$. During training, we keep the parameters of the image encoder, LLM, and text encoder frozen. We use AdamW to optimize the model with a learning rate of $1e^{-5}$ and the strategy of mixed precision.
Method NExT-QA Causal-VidQA Intent-QA
Des. Tem. Cau. All Des. Exp. Pre. Cou. All CW. CH. Tem. All
Co-Mem Gao et al. (2018) 54.4 50.0 45.9 48.5 64.1 62.8 31.4 32.6 47.7 47.7 54.9 39.1 46.8
HCRN Le et al. (2020) 54.0 49.3 47.1 48.9 56.4 61.6 32.6 32.7 48.1 - - - -
HME Fan et al. (2019) 57.4 48.9 46.8 49.2 63.4 61.5 28.9 30.9 46.2 46.1 54.3 40.8 46.2
B2A Park et al. (2021) 58.3 49.0 47.4 49.6 66.2 62.9 31.2 35.2 49.1 - - - -
HGA Jiang and Han (2020) 57.8 49.1 48.1 50.0 65.7 63.5 32.2 34.3 48.9 44.9 51.0 39.6 44.6
IGV Li et al. (2022c) 59.6 51.7 48.6 51.3 65.9 62.1 35.0 31.2 48.6 - - - -
HQGA Xiao et al. (2022a) 59.4 52.3 49.0 51.8 - - - - - 48.2 54.3 41.7 47.7
MCR Zang et al. (2023) 62.3 52.0 49.2 52.4 67.5 65.6 37.8 33.4 51.1 - - - -
VGT Xiao et al. (2022b) 67.3 54.5 52.8 55.7 70.8 70.3 38.4 42.0 55.4 51.4 56.0 47.6 51.3
CaVIR Li et al. (2023b) - - - - - - - - - 58.4 65.5 50.5 57.6
TranSTR Li et al. (2023e) 70.0 60.2 59.7 61.5 73.6 75.8 48.9 50.3 62.2 - - - -
SeViLA Yu et al. (2023) - - - 71.5 - - - - - - - - -
InstructBLIP Dai et al. (2023) 79.8 70.5 71.5 72.5 79.5 81.4 64.7 56.8 70.6 73.0 70.2 68.8 71.5
GCG (Ours) 80.7 72.6 74.2 74.6 80.7 82.3 66.5 59.1 72.1 75.0 71.9 69.2 73.1
Table 1: Accuracy (%) on NExT-QA, Causal-VidQA, and Intent-QA. Des, Tem, and Cau denote question types of Descriptive, Temporal, and Causal in NExT-QA. Des, Exp, Pre, and Cou denote question types of Description, Explanation, Prediction, and Counterfactual in Causal-VidQA. CW, CH, and Tem denote question types of Causal Why, Causal How, and Temporal in Intent-QA.
Method MSVD MSRVTT ActivityNet
MHN Peng et al. (2022) 40.4 38.6 -
VQA-T Yang et al. (2021) 46.3 41.5 38.9
TG-VQA Li et al. (2023a) 52.5 46.3 48.3
MuLTI-L Xu et al. (2023b) 54.7 47.8 -
FrozenBiLM Yang et al. (2022) 54.8 47.0 43.2
UMT-L Li et al. (2023d) 55.2 47.1 47.9
InternVideo Wang et al. (2022) 55.5 47.1 -
HiTea Ye et al. (2022) 55.6 45.9 46.4
mPLUG2 Xu et al. (2023a) 58.1 48.0 -
COSA-L Chen et al. (2023b) 58.6 48.8 49.2
VALOR-L Chen et al. (2023a) 60.0 49.2 48.6
InstructBLIP Dai et al. (2023) 59.7 47.0 46.3
GCG (Ours) 61.7 49.5 49.9
Table 2: Accuracy (%) on open-ended VideoQA datasets.
Settings NExT-QA Intent-QA MSVD-QA MSRVTT-QA
Baseline 72.5 71.5 59.7 47.0
Pseudo-labeled (Q) 72.3(-0.2) 70.3(-1.2) 59.9(+0.2) 47.5(+0.5)
Pseudo-labeled (E) 79.1(+6.6) 78.3(+6.8) 64.3(+4.6) 51.8(+4.8)
Table 3: As a preliminary step, we analyze the performance upper bound with weakly labeled keyframes as visual inputs.
5.3 Results and Analysis
The experimental results in Table 1 and 2 show the superiority of our weakly-supervised GCG compared to existing approaches. For the multi-choice setting, we achieve an accuracy of 74.6%, 72.1%, and 73.1% in NExT-QA, Causal-VidQA, and Intent-QA respectively. For the open-ended setting, we achieve an accuracy of 61.7%, 49.5%, and 49.9% in MSVD-QA, MSRVTT-QA, and ActivityNet-QA respectively. To ensure a fair comparison, we also apply the same settings to get the results for the vanilla InstructBLIP on these datasets as baselines, with $K=4$ frames uniformly sampled from the $T=32$ frames as visual inputs. Although the baseline InstructBLIP performed fair on these datasets, our proposed GCG showed a significant improvement, particularly in questions that require complex causal-temporal reasoning (+2.1% and +2.7% for Tem and Cau in NExT-QA, +2.3% for Cou in Causal-VidQA). Moreover, despite not having undergone video-text pretraining, our method still surpasses those pre-trained VideoQA models (e.g., HiTea, VALOR) on MSVD-QA, MSRVTT-QA, and ActivityNet-QA. We also observe that the improvements on ActivityNet-QA (+3.6%) are larger than the MSVD-QA (+2%) and MSRVTT-QA (+2.5%). This can be attributed to the average video length of ActivityNet-QA being 180 seconds, which is much longer than MSVD-QA (10s) and MSRVTT-QA (15s), demonstrating the necessity of discovering question-critical moments.
Settings NExT-QA MSVD-QA
Des. Tem. Cau. All
Baseline 79.8 70.5 71.5 72.5 59.7
$\mathcal{L}_{vqa}$ 80.3 69.9 72.3 72.9 59.5
$\mathcal{L}_{vqa}+\mathcal{L}_{reg}$ 79.3 70.5 73.5 73.6 60.4
$\mathcal{L}_{vqa}+\mathcal{L}_{con}$ 80.9 71.4 73.3 74.0 60.7
$\mathcal{L}_{vqa}+\mathcal{L}_{reg}+\mathcal{L}_{con}$ 80.7 72.6 74.2 74.6 61.7
Table 4: Ablation studies on loss components of GCG.
Figure 4: Qualitative results on NExT-QA test set. The frames selected by our method are highlighted in blue dashed lines. The ground truth answers are in green. We also display the weakly grounded similarity scores of each frame.
Figure 5: Ablation studies on hyperparameters $T$ and $\sigma$.
5.4 Pseudo-labels Analysis.
We further explore the performance with different frames as visual inputs to verify the pseudo-labeled $w$: Pseudo-labeled (E) means we choose the $K$ frames whose indexes correspond to $w\in\mathbb{R}^{K}$ as visual inputs (detailed in Section 4.2). Pseudo-labeled (Q) is obtained the same way as Pseudo-labeled (E) but with the pure question for similarity computation. Baseline means we uniformly sample $K$ frames as visual inputs.
Table 3 shows that Pseudo-labeled (E) exhibits significantly improved performance, particularly in benchmarks featuring longer videos and more complicated questions like NExT-QA and Intent-QA. This verifies the effectiveness of using the event description to provide pseudo-labels as weak supervision. Moreover, Pseudo-labeled (E) performs much better than Pseudo-labeled (Q). This can be explained from two perspectives: (1) The event descriptions include the contents of the answers needed to be grounded, filling the semantic gap between the pure question and the answer. (2) The CLIP models are mostly pre-trained on images and declarative texts, therefore the declarative event descriptions are more suitable for similarity computation to decide keyframes.
5.5 Ablation Studies
We investigate the role of our weakly-supervised framework with different variants and hyperparameters of GCG.
Loss components in GCG. We exhaust the combination of different loss components in GCG. Table 4 shows the results:
  • $\mathcal{L}_{vqa}$ solely hardly outperforms the baseline, because the Gaussian generator can not identify the causal scene without supervision of question-critical moments. This reflects our motivation in using the event descriptions to generate pseudo labels as weak supervision.
  • $\mathcal{L}_{vqa}+\mathcal{L}_{reg}$ and $\mathcal{L}_{vqa}+\mathcal{L}_{con}$ match equally in accuracy that consistently surpasses baseline and $\mathcal{L}_{vqa}$. $\mathcal{L}_{reg}$ is responsible for regularizing the indexes of peaks in $p$ to approximate the timestamps of weakly grounded $w$, and $\mathcal{L}_{con}$ ensures the maximization between the selected moments and event descriptions.
  • $\mathcal{L}_{vqa}+\mathcal{L}_{reg}+\mathcal{L}_{con}$ is the complete GCG, which further boosts the performance significantly in all cases, showing that $\mathcal{L}_{reg}$ and $\mathcal{L}_{con}$ contribute in different aspects and their benefits are mutually reinforcing.
Number of frames $T$ and Gaussian widths $\sigma$. We evaluate how the number of overall sampled frames $T$ and the widths $\sigma$ of Gaussian functions will influence the performance. Figure 5 (a) indicates that performance improves as more frames are included, however, beyond a certain threshold ($T=48$), there is a performance drop. This suggests that too many frames may introduce redundancy and noise, while too few frames miss important information. As for $\sigma$ in Equation 4, it essentially decides the width and the degree of dispersion of the Gaussian distribution $g$. A larger $\sigma$ generates more dispersed $g$ with a more exploratory $p$ and vice versa. In Figure 5 (b), we vary $\sigma$ from 0.1 to 0.7 and observe that the performance fluctuates in a range of [73.9, 74.6] for NExT-QA and [61.4, 61.7] for MSVD-QA. The performance of NExT-QA is more sensitive to $\sigma$, and we argue that this is because the average video length of NExT-QA (44s) is much longer than MSVD-QA (10s).
5.6 Qualitative results.
We also present qualitative results in Figure 4 to gain a clearer insight into our method, along with the frames identified by GCG (bounded with blue dashed lines) and the weakly grounded similarities computed by the event description. Both cases show semantic correspondence between the question and the selected moments. For instance, in the upper case, the frames identified by GCG precisely correspond to the event [The baby rolled the roller on the toy when she reached it] with relatively higher similarity scores, demonstrating the strong ability of our GCG to localize video moments crucial for answering the question. By leveraging the information from this localized segment, the LMM can successfully arrive at the correct answer [ roll the roller on the toy]. In contrast, without our GCG, the presence of massive redundancy in the video overwhelms the reasoning process and leads to a false prediction of [laugh loudly].
6 Conclusion
In this paper, we have studied the problem of discovering question-critical moments in videos when adapting LMMs for VideoQA tasks. To address the shortcomings of the uniform sampling strategy and the absence of human annotations for question-critical timestamps in VideoQA datasets, we introduce a weakly-supervised framework to force the LMMs to reason out the answers by grounding question-critical moments. We propose the Gaussian-based Contrastive Grounding, a flexible and lightweight method to dynamically select keyframes with end-to-end training. Through a series of experiments and analyses, we have demonstrated the effectiveness of our approach in various challenging VideoQA tasks, particularly excelling in causal-temporal reasoning. For future work, we intend to focus on mitigating the biases and superficial correlations observed in previous studies, and further enhance the reasoning abilities of current models.